OpenStudy (anonymous):

Use the Product Rule to differentiate the function. Simplify your answer. \[g(v) = (v-\sqrt{v})(v^2+\sqrt{v})\]

OpenStudy (callisto):

\[f(v) = (v-\sqrt{v}), h(v) = (v^2+\sqrt{v})\] What are f'(v) and h'(v)?

OpenStudy (anonymous):

f'(x)=1-1/2v^-1/2 h'(x) = 2v+1/2v^-1/2 ???

OpenStudy (callisto):

Next, \[g'(v) = f'(v)h(v) + h'(v)f(v) = (1-\frac{1}{2\sqrt{v}})(2v+\frac{1}{2\sqrt{v}})\] Expand and simplify it?!

OpenStudy (anonymous):

i think this is the answer: \[3v^2-\frac{ 3 }{ 2 }\sqrt{v}-\frac{ 5 }{ 2 }v ^{3/2}-1\]

OpenStudy (callisto):

How do you get it? D:

OpenStudy (callisto):

Oh! My apology!! I did the wrong substitution!!

OpenStudy (callisto):

\[g'(v) = f'(v)h(v) + h'(v)f(v) = (1-\frac{1}{2\sqrt{v}}) (v^2 + \sqrt{v}) + (v-\sqrt{v})(2v+\frac{1}{2\sqrt{v}})\] This is the right one. I'm sorry!!!

OpenStudy (anonymous):

it's okay, i was reading my textbook and saw the formula for Product Rule

OpenStudy (anonymous):

I just have problems dealing with the fraction exponents... :(

OpenStudy (callisto):

Still have problems?!

OpenStudy (anonymous):

...

OpenStudy (callisto):

What do those dots mean?

OpenStudy (anonymous):

I still got \[g'(x)= 3v^2-\frac{ 5 }{ 2 }v ^{3/2}+\frac{ 3 }{ 2 }\sqrt{v}-1\]

OpenStudy (callisto):

Which is correct?!

OpenStudy (anonymous):

Is it correct? The question mark confused me.

OpenStudy (callisto):

That's what I get.

OpenStudy (anonymous):

Is there an easier way to deal with fraction exponents? sometimes i make mistakes because of exponents

OpenStudy (anonymous):

by the way, thanks for helping me out with my derivative questions! Thank you so much! :)

OpenStudy (callisto):

Sometimes, factoring may help, it depends. I guess just practise more. You're welcome :)